# Overdetermined Steklov eigenvalue problems on compact surfaces

### v1

@@ 4.1. Riemannian surfaces
Let $(\Omega,g_{\Omega})$ be a connected compact surface with $k$ boundary components, $k\geq 2$. Then $\Omega$ is solvable for the overdetermined problem (OS-I) if and only if $\Omega\sim\mathbb{A}_{L}$ for some $L\geq L_{0}$, where $\mathbb{A}_{L}$ and $L_{0}$ are defined in Sect. 1.
First, let us check the solvability of the cylinder $\mathbb{A}_{L}$. It is a special case of [GP17]*Example 1.3.3 (originally from [CESG11]*Lemma 6.1) and shows that both Proposition 3.3 and Theorem 3.1 no longer hold if the simply connected condition is removed.
@@
The boundary $\partial\mathbb{A}_{L}$ has two connected components $\Gamma_{\pm}:=\{\pm L\}\times S^{1}$.
By the method of separation of variables, the Steklov spectrum of $\mathbb{A}_{L}$ is given by
$0,\quad 1/L,\quad k\tanh(kL),\quad k\coth(kL)\quad(k\geq 1),$
@@
Since $\tanh(kL)\leq\coth(kL)$ for any $k\geq 1$ and $\tanh(\cdot)$ is strictly increasing, we have $\sigma_{1}=1/L$ or $\sigma_{1}=\tanh L$. By the monotonicity of $1/L$ and $\tanh L$ with respect to $L$ and the asymptotic behaviors as $L$ tends to 0 and $\infty$, there is a unique constant $L_{0}>0$ such that $1/L_{0}=\tanh L_{0}$.
(1) When $L\geq L_{0}$, we have $1/L\leq\tanh L$. Hence, $\sigma_{1}=1/L$. Moreover, $\sigma_{1}$ is simple when $L>L_{0}$ while $\sigma_{1}$ has multiplicity $3$ when $L=L_{0}$. In this case, $u_{1}=t$ solves (OS-I), since $\nabla u_{1}=\partial_{t}$ and then $|\nabla u_{1}|\equiv 1$ on $\bar{\Omega}$.
Moreover, $u_{1}\equiv L$ on $\Gamma_{+}$ while $u_{1}\equiv-L$ on $\Gamma_{-}$, i.e., $u_{1}$ is constant on each connected component of the boundary respectively but is not constant on the whole boundary. Meanwhile, $\frac{1}{2}\frac{\partial}{\partial\nu}\log|\nabla u_{1}|^{2}=0=-\kappa$, which coincides with Claim 3 in the proof of Proposition 3.3, but Proposition 3.3 itself fails.
(2) When $L<L_{0}$, we have $\tanh L<1/L$. Hence, $\sigma_{1}=\tanh L$, which has multiplicity $2$. In this case, any first Steklov eigenfunction can be written as
@@
$|\nabla u_{1}|^{2}=(A\cos\theta+B\sin\theta)^{2}\sinh^{2}L+(B\cos\theta-A\sin\theta)^{2}\cosh^{2}L=:Q(\theta)$
@@
$A^{2}\sinh^{2}L+B^{2}\cosh^{2}L=Q(0)=Q(\pi/2)=B^{2}\sinh^{2}L+A^{2}\cosh^{2}L$
@@
The cylinder $\mathbb{A}_{L}$ is solvable for the overdetermined problem (OS-I) if and only if $L\geq L_{0}$, where $L_{0}$ is the unique positive solution of the equation $\tanh x=1/x$.
@@
Assume $\partial\Omega$ has $k$ connected components, denoted by $\Gamma_{1},\cdots,\Gamma_{k}(k\geq 2)$. Let $u$ be a non-trivial solution of (OS-I) and $f=\log|\nabla u|^{2}$. For each $i$, we have the following assertions.
@@
$\frac{1}{2}\frac{\partial f}{\partial\nu}=-\kappa\mbox{ on }\Gamma_{i}.$
@@
$\frac{1}{2}\frac{\partial f}{\partial\nu}=\sigma_{1}-\kappa\mbox{ on }\Gamma_{i}.$
@@
Assume the solvability of (OS-I). By Lemma 3.2, Proposition 4.4 and Gauss–Bonnet formula, we have
$\int_{\Omega}K$ $=\int_{\Omega}\frac{1}{2}\Delta f=\int_{\partial\Omega}\frac{1}{2}\frac{\partial f}{\partial\nu}=\sum_{i=1}^{k}\int_{\partial\Gamma_{i}}\delta_{i}\sigma_{1}-\int_{\partial\Omega}\kappa$
$=\sum_{i=1}^{k}\sigma_{1}\delta_{i}L(\partial\Gamma_{i})+\int_{\Omega}K-2\pi\chi(\Omega),$
@@
(4.1) $\sigma_{1}\sum_{i=1}^{k}\delta_{i}L(\partial\Omega)=2\pi\chi(\Omega)=2\pi(2-2\gamma-k),$
@@
Since the left-hand side of (4.1) is nonnegative, (4.1) cannot hold if $\gamma\geq 1$ or $k\geq 3$. Thus, we must have $\gamma=0$ and $k\leq 2$. We rule out the case $\gamma=0,k=1$ since $\Omega$ would be simply connected.
When $\gamma=0$ and $k=2$, $\Omega$ is topologically an annulus. The right-hand side of (4.1) equals to 0, which forces $\delta_{1}=\delta_{2}=0$. By Proposition 4.4, $u$ must be constant on each $\Gamma_{i}$, denoted by $c_{i}$. Then we deduce $c_{i}^{2}=c^{2}/\sigma_{1}^{2}$ on $\Gamma_{i}$ from (3.3) and (3.5) with $u_{y}(0,y)$. On the other hand, (2.6) implies $c_{1}L(\Gamma_{1})+c_{2}L(\Gamma_{2})=0$, hence, $c_{1}=-c_{2}$ and $L(\Gamma_{1})=L(\Gamma_{2})$.
@@
(4.2) $\Phi\colon\Omega\to\mathbb{A}_{L}.$
We now show that $\Phi$, up to scaling, is an isometry on the boundary. Without loss of generality, we assume $\Phi(\Gamma_{1})=\Gamma_{-},\Phi(\Gamma_{2})=\Gamma_{+},L(\Gamma_{1})=L(\Gamma_{2})=L(\Gamma_{-})=L(\Gamma_{+})$.
Recalling Example 4.2, let $\ell(t)=t$ be the first eigenfunction of $\mathbb{A}_{L}$ which is linear in $t$ and normalized by $\|\ell\|_{L^{2}(\partial\mathbb{A}_{L})}=1$. Pulling $\ell$ back to $\Omega$, we obtain $\|\ell\circ\Phi\|_{L^{2}(\partial\Omega)}=1$.
The restriction of $\ell$ to $\Gamma_{-}$ and $\Gamma_{+}$ is constants, denoted by $l_{-}$ and $l_{+}$, respectively. Clearly, $l_{-}+l_{+}=0$, then we have
(4.3) $\int_{\partial\Omega}\ell\circ\Phi=l_{-}L(\Gamma_{1})+l_{+}L(\Gamma_{2})=0,$
Therefore, we may use $\ell\circ\Phi$ as a test function in the variational characterization of $\sigma_{1}(\Omega)$ to obtain
(4.4) $\sigma_{1}(\Omega)\leq E(\ell\circ\Phi)=E(\ell)=\sigma_{1}(\mathbb{A}_{L}).$
Conversely, consider $\Phi^{-1}\colon\mathbb{A}_{L}\to\Omega$ and the eigenfunction $u$ of $\Omega$ solving (OS-I). Since it has been proved that $u|_{\Gamma_{i}}=c_{i}(i=1,2)$ and $c_{1}+c_{2}=0$, then by the same argument above, we may use $u\circ\Phi^{-1}$ as the test function to estimate $\sigma_{1}(\mathbb{A}_{L})$, i.e.,
(4.5) $\sigma_{1}(\mathbb{A}_{L})\leq E(u\circ\Phi^{-1})=E(u)=\sigma_{1}(\Omega).$
Hence, $\sigma_{1}(\mathbb{A}_{L})=\sigma_{1}(\Omega)=\sigma_{1}$, and $\ell\circ\Phi$ is a first eigenfunction of $\Omega$. If we write the metric of $\Omega$ as $\rho^{2}(\mathrm{d}t^{2}+\mathrm{d}\theta^{2})$, then $\frac{\partial\ell}{\partial\nu}=\sigma_{1}\ell$ on $\partial\Omega$ where $\nu=-\rho^{-1}\partial_{t}$ on $\Gamma_{1}$ and $\nu=\rho^{-1}\partial_{t}$ on $\Gamma_{2}$. Since both $\ell$ and $\partial_{t}\ell$ are constant on each boundary component, so is $\rho$. Finally, we conclude $\rho\equiv 1$ on the boundary since we have assumed $\Gamma_{+}$ and $\Gamma_{1}$ have the same length. This completes the whole proof. ∎
@@ 4.2. Domains in space forms
Suppose $\Omega\subset\mathbb{M}^{2}_{K}$ is solvable for (OS-I), then $\Omega\sim\mathbb{A}_{L}$ for some $L$ by Theorem 4.1. We write the metric $g$ of $\Omega$ as $g=\rho^{2}g_{0}$ with $g_{0}=\mathrm{d}t^{2}+\mathrm{d}\theta^{2}$, then
@@
(1) $\Omega\subset\mathbb{R}^{2}(K=0)$: The equation (4.6) becomes $\Delta_{0}w=0$, i.e., $w$ is harmonic. Since $w|_{\partial\Omega}$ is constant, we deduce that both $w$ and $\rho=e^{w}$ are constant. Thus, $\Omega$ is isometric to $\mathbb{A}_{L}$ after rescaling. It is a contradiction, since each connected component of $\partial\mathbb{A}_{L}$ is a geodesic, but each complete geodesic in $\mathbb{R}^{2}$ is a straight line and it is non-compact.
(2) $\Omega\subset\mathbb{H}^{2}(K=-1)$: The equation (4.6) becomes $\Delta_{0}w=e^{2w}$. Since $w|_{\partial\Omega}$ is constant, this semi-linear elliptic equation has at most one solution by the maximum principle. If such a solution exists, it is independent of $\theta$ and is even in $t$, i.e., $w=w(t)$ and w(t)=w(t)w(t)=w(-t); so is $\rho=e^{w}$. Consequently, the geodesic curvature of each boundary component is constant. Since the boundary components have constant geodesic curvature and equal length, both $\Gamma_{1}$ and $\Gamma_{2}$ are closed geodesic circles of the same radius. However, such circles cannot be nested and do not bound a compact domain. Hence, it is impossible that $\Omega\sim\mathbb{A}_{L}$.
@@
A multiply connected domain $\Omega\subset\mathbb{S}^{2}$ with boundary $\partial\Omega$ is solvable for (OS-I) if and only if $\Omega\sim A_{\mathrm{sph}}(R)$ for certain $R\in[R_{0},\pi/2)$.
@@
This implies $A_{\mathrm{sph}}(R)\sim\mathbb{A}_{L}$ with $R=2\arctan\tanh(L/2)$. Moreover, $t=0\iff r=0$ while $L_{0}\leq L<+\infty\iff R_{0}\leq R<\pi/2$.
Hence, $\Omega$ is solvable for (OS-I) if and only if $\Omega\sim\mathbb{A}_{L}\sim A_{\mathrm{sph}}(R)$ for certain $R\in[R_{0},\pi/2)$ by Theorem 4.1. ∎

### v2

@@ 4.1. Riemannian surfaces
Let $(\Omega,g_{\Omega})$ be a multiply connected compact surface. Then $\Omega$ is solvable for the overdetermined problem (OS-I) if and only if $\Omega\sim\mathbb{A}_{T}$ for some $T\geq T_{1}$, where $\mathbb{A}_{T}$ and $T_{1}$ are defined in Sect. 1.
First, let us check the solvability of the cylinder $\mathbb{A}_{T}$. It is a special case of [GP17]*Example 1.3.3 (originally from [CESG11]*Lemma 6.1) and shows that both Proposition 3.3 and Theorem 3.1 no longer hold if the simply connected condition is removed.
@@
The boundary $\partial\mathbb{A}_{T}$ has two connected components $\Gamma_{\pm}\coloneq\{\pm T\}\times S^{1}$.
By the method of separation of variables, the Steklov spectrum of $\mathbb{A}_{T}$ is given by
$0,\quad 1/T,\quad k\tanh(kT),\quad k\coth(kT)\quad(k\geq 1),$
@@
Since $\tanh(kT)\leq\coth(kT)$ for any $k\geq 1$ and $\tanh(\cdot)$ is strictly increasing, we have $\sigma_{1}=1/T$ or $\sigma_{1}=\tanh T$. By the monotonicity of $1/T$ and $\tanh T$ with respect to $T$ and the asymptotic behaviors as $T$ tends to 0 and $\infty$, there is a unique constant $T_{1}>0$ such that $1/T_{1}=\tanh T_{1}$.
(1) When $T\geq T_{1}$, we have $1/T\leq\tanh T$. Hence, $\sigma_{1}=1/T$. Moreover, $\sigma_{1}$ is simple when $T>T_{1}$ while $\sigma_{1}$ has multiplicity $3$ when $T=T_{1}$. In this case, $u_{1}=t$ solves (OS-I), since $\nabla u_{1}=\partial_{t}$ and then $|\nabla u_{1}|\equiv 1$ on $\mathbb{A}_{T}$.
Moreover, $u_{1}\equiv T$ on $\Gamma_{+}$ while $u_{1}\equiv-T$ on $\Gamma_{-}$, i.e., $u_{1}$ is constant on each connected component of the boundary respectively but is not constant on the whole boundary. Meanwhile, $\frac{\partial}{\partial\nu}\log|\nabla u_{1}|=0=-\kappa$, which coincides with Claim 3 in the proof of Proposition 3.3, but Proposition 3.3 itself fails.
(2) When $T<T_{1}$, we have $\tanh T<1/T$. Hence, $\sigma_{1}=\tanh T$, which has multiplicity $2$. In this case, any first Steklov eigenfunction can be written as
@@
$|\nabla u_{1}|^{2}=(A\cos\theta+B\sin\theta)^{2}\sinh^{2}T+(B\cos\theta-A\sin\theta)^{2}\cosh^{2}T\eqcolon Q(\theta)$
@@
$A^{2}\sinh^{2}T+B^{2}\cosh^{2}T=Q(0)=Q(\pi/2)=B^{2}\sinh^{2}T+A^{2}\cosh^{2}T$
@@
The cylinder $\mathbb{A}_{T}$ is solvable for the overdetermined problem (OS-I) if and only if $T\geq T_{1}$, where $T_{1}$ is the unique positive solution of the equation $\tanh x=1/x$.
@@
Let $(\Omega,g_{\Omega})$ be a connected orientable compact surface. Assume $\partial\Omega$ has $b$ connected components, denoted by $\Gamma_{1},\cdots,\Gamma_{b}(b\geq 1)$. Let $u$ be a nontrivial solution of (OS-I) and $f=\log|\nabla u|$. For each $i$, we have the following assertions.
@@
$\frac{\partial f}{\partial\nu}=-\kappa\text{ on }\Gamma_{i}.$
@@
$\frac{\partial f}{\partial\nu}=\sigma_{1}-\kappa\text{ on }\Gamma_{i}.$
Moreover, Assertion (1) does not occur when $b=1$.
@@
(I) Orientable Case. Assume the solvability of (OS-I) for an orientable surface $\Omega$. By Lemma 3.2, Proposition 4.4 and Gauss–Bonnet formula, we have
$\int_{\Omega}K$ $=\int_{\Omega}\Delta f=\int_{\partial\Omega}\frac{\partial f}{\partial\nu}=\sum_{i=1}^{b}\int_{\Gamma_{i}}\delta_{i}\sigma_{1}-\int_{\partial\Omega}\kappa$
$=\sum_{i=1}^{b}\sigma_{1}\delta_{i}L(\Gamma_{i})+\int_{\Omega}K-2\pi\chi(\Omega),$
@@
(4.1) $\sigma_{1}\sum_{i=1}^{b}\delta_{i}L(\Gamma_{i})=2\pi\chi(\Omega)=2\pi(2-2\gamma-b),$
@@
Since the left-hand side of (4.1) is nonnegative, (4.1) cannot hold if $\gamma\geq 1$ or $b\geq 3$. Thus, we must have $\gamma=0$ and $b\leq 2$. We rule out the case $\gamma=0,b=1$ since $\Omega$ would be simply connected.
When $\gamma=0$ and $b=2$, $\Omega$ is topologically an annulus. The right-hand side of (4.1) is equal to 0, which forces $\delta_{1}=\delta_{2}=0$. By Proposition 4.4, $u$ must be constant on each $\Gamma_{i}$, denoted by $c_{i}$. Then we deduce $c_{i}^{2}=c^{2}/\sigma_{1}^{2}$ on $\Gamma_{i}$ from (3.2) and (3.4) with $u_{y}(0,y)$. On the other hand, $\int_{\partial\Omega}u=0$ implies $c_{1}L(\Gamma_{1})+c_{2}L(\Gamma_{2})=0$, hence, $c_{1}=-c_{2}$ and $L(\Gamma_{1})=L(\Gamma_{2})$.
Now by Proposition 4.5 (see below) and Example 4.2, we obtain that $\Omega$ is $\sigma$-homothetic to $\mathbb{A}_{T}$ for some $T\geq T_{1}$.
(II) Non-orientable case. Assume that $(\Omega,g)$ is non-orientable. We consider its orientable double cover $\tilde{\Omega}$ with the $2:1$ projection $\pi\colon\tilde{\Omega}\to\Omega$ and the metric $\tilde{g}=\pi^{\ast}g$. Let $\tilde{u}=\pi^{\ast}u=u\circ\pi$. Clearly, if $u$ has no critical points in $\Omega$, then $\tilde{u}$ has no critical points in $\tilde{\Omega}$. If $u$ solves (OS-I) for $\Omega$, then $\tilde{u}$ solves (OS-I) for $\tilde{\Omega}$. Although $\sigma_{1}=\sigma_{1}(\Omega)$ may not be the first nonzero Steklov eigenvalue of $\tilde{\Omega}$, Lemma 3.2 and Proposition 4.4 are still valid. Similar to the proof of Theorem 4.1, we obtain $\tilde{\Omega}$ is topologically an annulus, and $\tilde{u}$ takes opposite constant values on the two boundary components. This is a contradiction.
This completes the whole proof. ∎
We now prove the following proposition used in the proof of Theorem 4.1.
Proposition 4.5.
Let $\Omega$ be topologically an annulus equipped with a smooth metric $g_{\Omega}$. If $\Omega$ has a Steklov eigenfunction $u$ such that the restriction of $u$ to the two boundary components are constants of opposite sign, then there exists a $\sigma$-homothety $\Phi\colon\Omega\to\mathbb{A}_{T}$ for some $T>0$, and $u$ is $\Phi^{\ast}t$ times a nonzero constant.
Proof.
@@
$\Phi\colon\Omega\to\mathbb{A}_{T}$
for some $T>0$. Without loss of generality, we assume $\Phi(\Gamma_{1})=\Gamma_{-}$ and $\Phi(\Gamma_{2})=\Gamma_{+}$.
By the conformal map $\Phi$, we can parametrize $\Omega$ by $(t,\theta)\in[-T,T]\times S^{1}$ and write the metric as $g_{\Omega}=\rho^{2}(t,\theta)(\mathrm{d}t^{2}+\mathrm{d}\theta^{2})$.
Suppose $u$ is a Steklov eigenfunction such that $u|_{\Gamma_{1}}=u|_{\Gamma_{2}}=c$, where $c$ is a nonzero constant. Then $u$ is the unique solution of the following boundary value problem
$\begin{cases}u_{tt}+u_{\theta\theta}=0,\\ u(-T,\theta)=-c,u(T,\theta)=c.\end{cases}$
We can directly verify that the linear function $u=ct/T$ is the unique solution.
Since $u$ is a nonconstant Steklov eigenfunction, we have $\frac{\partial u}{\partial\nu}=\sigma u$ on $\partial\Omega$, where $\sigma>0$, $\nu=-\rho^{-1}\partial_{t}$ on $\Gamma_{1}$, and $\nu=\rho^{-1}\partial_{t}$ on $\Gamma_{2}$. This implies
$\rho^{-1}(-T,\theta)\frac{c}{T}=-c\sigma,\quad\rho^{-1}(T,\theta)\frac{c}{T}=c\sigma.$
Hence, $\rho=1/(\sigma T)$ is constant on $\partial\Omega$ since $c\neq 0$, which implies $\Phi$ is a $\sigma$-homothety. ∎
@@ 4.2. Domains in space forms
Suppose $\Omega\subset\mathbb{M}^{2}_{K}$ is solvable for (OS-I), then $\Omega\sim\mathbb{A}_{T}$ for some $T$ by Theorem 4.1. We write the metric $g$ of $\Omega$ as $g=\rho^{2}g_{0}$ with $g_{0}=\mathrm{d}t^{2}+\mathrm{d}\theta^{2}$, then
@@
(1) $\Omega\subset\mathbb{R}^{2}~(K=0)$: The equation (4.2) becomes $\Delta_{0}w=0$, i.e., $w$ is harmonic. Since $w|_{\partial\Omega}$ is constant, we deduce that both $w$ and $\rho=e^{w}$ are constant. Thus, $\Omega$ is isometric to $\mathbb{A}_{T}$ after rescaling. It is a contradiction, since each connected component of $\partial\mathbb{A}_{T}$ is a geodesic, but each complete geodesic in $\mathbb{R}^{2}$ is a straight line and it is non-compact.
(2) $\Omega\subset\mathbb{H}^{2}~(K=-1)$: The equation (4.2) becomes $\Delta_{0}w=e^{2w}$. Since $w|_{\partial\Omega}$ is constant, this semi-linear elliptic equation has at most one solution by the maximum principle. If such a solution exists, it is independent of $\theta$ and is even in $t$, i.e., $w=w(t)$ and w(t)=w(t)w(t)=w(-t); so is $\rho=e^{w}$. Consequently, the geodesic curvature of each boundary component is constant. Since the boundary components have constant geodesic curvature and equal length, both $\Gamma_{1}$ and $\Gamma_{2}$ are closed geodesic circles of the same radius. However, such circles cannot be nested and do not bound a compact domain. Hence, it is impossible that $\Omega\sim\mathbb{A}_{T}$.
@@
A multiply connected domain $\Omega\subset\mathbb{S}^{2}$ with boundary $\partial\Omega$ is solvable for (OS-I) if and only if $\Omega\sim A_{\mathrm{sph}}(R)$ for some $R\in[R_{1},\pi/2)$. Moreover, $\Omega$ is symmetric with respect to a certain great circle of $\mathbb{S}^{2}$.
@@
This implies $A_{\mathrm{sph}}(R)\sim\mathbb{A}_{T}$ with $R=2\arctan\tanh(T/2)$. Moreover, $t=0\iff r=0$ while $T_{1}\leq T<+\infty\iff R_{1}\leq R<\pi/2$.
Hence, $\Omega$ is solvable for (OS-I) if and only if $\Omega\sim\mathbb{A}_{T}\sim A_{\mathrm{sph}}(R)$ for some $R\in[R_{1},\pi/2)$ by Theorem 4.1.
When $K=1$, the equation (4.2) becomes $\Delta_{0}w=-e^{2w}$. The boundary condition is $w(T,\cdot)=w(-T,\cdot)=c$. By using the moving plane method (cf. the proof of Theorem 2.1 in [GNN79]), we can prove that the solution $w$ of this boundary value problem is symmetric with respect to $t$, and $\frac{\partial w}{\partial t}>0$ for $t\in(-T,0)$ while $\frac{\partial w}{\partial t}<0$ for $t\in(0,T)$. Therefore, $\rho=e^{w}$ is also symmetric with respect to $t$ and $\rho_{t}(0,\cdot)=0$. It follows that the image of $\{0\}\times S^{1}$ under a $\sigma$-homothety $\Phi\colon\mathbb{A}_{T}\to\Omega$ is a closed geodesic of $\mathbb{S}^{2}$. Hence, $\Omega$ is symmetric with respect to a certain great circle of $\mathbb{S}^{2}$. ∎